# Bolt3D: Generating 3D Scenes in Seconds

### v1

@@ 5.4 Image VAEs generalize poorly to geometry
Rel.$\times$100$\downarrow$ $\delta_{1.01}$$\uparrow$ $\Delta uv(\text{px})$$\downarrow$
@@
We use metrics commonly used for depth estimation [39, 58]—relative error (AbsRel) between the target z-component of the pointmap $z$ and prediction $\hat{z}$, Rel. $=\left\lVert z-\hat{z}\right\rVert/z$, and the prediction threshold accuracy, $\delta_{\epsilon}=\operatorname{max}(\nicefrac{{\hat{z}}}{{z}},\nicefrac{{z}}{{\hat{z}}})<\epsilon$ with $\epsilon=1.01$ Additionally, we measure the re-projection error: the mean Euclidean distance on the image plane when re-projecting the point back to the source camera $\Delta uv=\frac{1}{N}\sum_{i}^{N}\left\lVert(u_{i},v_{i})-(\hat{u}_{i},\hat{v}_{i})\right\rVert_{2}$, where $u$ and $v$ are the $x$ and $y$ coordinates, respectively, in pixels on the image. We measure performance separately for synthetic and real data, both at $512\times 512$ resolution.
@@
Our geometry data is not metric, so the scale of the scene is arbitrary. Properly normalizing the data before inputting it to the VAE can significantly impact performance. We consider 3 methods for scaling data. We experiment with (1) scaling the scene such that mean scene depth is $1$, (2) scaling the scene such that the maximum coordinate value in any direction is $1$ and (3) applying a nonlinear contraction function (the sigmoid function) before inputting it to the encoder, and applying the inverse of that function after decoding. Our autoencoder uses (1).
@@ 6 Conclusion
We have presented Bolt3D, a fast feed-forward method that generates detailed 3D scenes in less than 7 seconds. To enable this capability, we propose a 3D scene representation that allows for denoising high-resolution 3D scenes using powerful 2D latent diffusion architectures. We create a large-scale 3D scene dataset to train Bolt3D, and demonstrate that it accurately models ambiguity, enabling high-quality 1-view reconstructions where regression-based methods fail. Our proposed feed-forward approach is capable of reconstructing a wide variety of scenes and reduces the cost of 3D generation by 300×300$\times$300 × compared to existing optimization-based methods, opening up opportunities for 3D content creation at scale.
@@ Acknowledgments
We would like to thank George Kopanas, Ben Poole, Sander Dieleman, Songyou Peng, Rundi Wu, Alex Trevithick, Daniel Duckworth, Dor Verbin, Linyi Jin, Hadi Alzayer, David Charatan, Jiapeng Tang and Akshay Krishnan for valuable discussions, insights and contributions. We also extend our gratitude to Shlomi Fruchter, Kevin Murphy, Mohammad Babaeizadeh, Han Zhang and Amir Hertz for training the base text-to-image latent diffusion model.
@@ 7.2 Ablations
Next, we evaluate performance at a lower resolution, closer to the training setup of baseline methods. In Tab. 8 we illustrate that at a lower $256\times 256$ resolution, our model still outperforms Depthsplat. Finally, we give Depthsplat an advantage by evaluating its performance when receiving wide field-of-view (fov) $256\times 448$ images. In this setting, the wider field-of-view provides more scene coverage by a factor of ×$\times 1.75$ and more cues for matching across different views. Only in this setting do we find that Depthsplat achieves similar performance to our method.
Rel. $\downarrow$ $\delta_{1.01}$uparrow $\Delta uv(\text{px})\downarrow$
@@
1-view RE10K Flash3D 256$\times$256 17.70 0.616 0.393
Ours 256$\times$256 21.62 0.804 0.202
3-view RE10K Depthsplat 256$\times$256 24.69 0.873 0.126
Ours 256$\times$256 27.39 0.916 0.103
2-view DL3DV Depthsplat 256$\times$448 18.09 0.549 0.323
Depthsplat 256$\times$256 16.16 0.467 0.388
Ours 256$\times$256 18.01 0.556 0.320
4-view DL3DV Depthsplat 256$\times$448 21.20 0.697 0.208
Depthsplat 256$\times$256 19.64 0.633 0.254
Ours 256$\times$256 21.16 0.695 0.231
6-view DL3DV Depthsplat 256$\times$448 21.93 0.730 0.184
Depthsplat 256$\times$256 20.64 0.680 0.225
Ours 256$\times$256 22.18 0.733 0.206
Table 8: Low-resolution 256$\times$256 comparisons. Our method outperforms competitors at low resolution when receiving the same input information. Only when DepthSplat [63] receives 1.75 times more input information than our method by ingesting wide-fov images, does its performance become similar to ours.
@@ 8.1 XYZ normalization
[…]
@@
[…]
@@
where $\left[R~{}|~{}T\right]_{\text{w2c}}$ is the world-to-camera transformation matrix. Because scenes are scaled such that the mean depth to the first camera is 1, we can think of $[0,0,1]^{\top}$ as the look-at point or center of the scene, so $d=\left\lVert\mathbf{x}_{\text{local}}-[0,0,1]^{\top}\right\rVert$ is the distance of the point to the center of the scene of the local camera. Thus, we compute:
@@
We use a U-Net with full 3D attention on all feature maps up to $32\times 32$, as in CAT3D [12]. Unlike CAT3D, our diffusion model is trained to model the joint distribution of latent appearance and geometry. To this end, we increase the number of channels in the input and output layers of CAT3D’s architecture by 8 to additionally accept geometry latents. The input to our network thus has 8-dimensions for the geometry latents, 8-dimensions for the image latent, 6-dimensions for the camera pose raymaps, and a 1-dimensional mask indicating which views are given as conditioning, yielding an input dimension of $64\!\times\!64\!\times\!23$. We train with the same optimization hyperparameters as [12], except we additionally finetune on $16$ input views with a lower learning rate of $51e-$ 5.
@@
We use a pre-trained and frozen image autoencoder similar to that of Stable Diffusion [42]. The geometry encoder has the same architecture, except we increase the channel dimension to additionally accept a 6-dimensional camera pose representation. The decoder is a transformer-based network. We patchify the $64\times 64\times 8$ latent with patch size $2$, thus using a token length $1024$. We use the ViT-B architecture hyperparameters: 12 layers with channel size $768$, with the fully-connected layer consisting of 2 dense layers with GeLU activation function and a hidden MLP dimension $3072$. The linear projection head projects each token to a $16\times 16$ patch. We optimize the parameters of the Autoencoder with the Adam [10] optimizer using constant learning rate $41e-$ 4, batch size $512$, Adam parameters $(\beta_{1},\beta_{2})=(0.0,0.99)$. We first train for 3M iterations at $256\times 256$ resolution, followed by fine-tuning at $512\times 512$ for 250k iterations. We use loss weight parameters $\lambda_{1}=3e-9$ and $\lambda_{2}=0.033$.
@@
We detail the Gaussian head architecture in Fig. 8. The Gaussian Head receives as input the 3-dimensional image, 3-dimensional pointmap, and 6-dimensional raymap encoding the camera pose. Each view is first passed through 3 residual convolutional blocks with the Swish activation function [38] and channel size $128$, followed by $4\times 4$ patchification to token dimension $128$ and full cross attention. We use 3 transformer layers with hidden dimension $128$ and $8$ heads, MLP dimension $512$. The tokens are then unpatchified to original resolution and $128$ channel dimension. Following that, there are another 3 residual convolutional blocks, and a final, unactivated $3\times 3$ convolutional layer that outputs channel size $11$ ($3$ for color, $3$ for size, $4$ for rotation and $1$ for opacity). The outputs are activated with an exponential function for scale and a sigmoid function for opacity. To facilitate accurate scale prediction, the size output by the network is then multiplied by the z-distance of the gaussian from the camera. The means of Gaussians are not predicted by the Gaussian head, as they are already available from the VAE. The Gaussian head is trained with $8$ input views, leading to a sequence length $131k$. We manage this computational complexity by using FlashAttention [8, 7] and rematerializing gradients on the dot product operation. For losses, we use L2 photometric loss with weight $\lambda=1$ and LPIPS loss weight $\lambda_{LPIPS}=0.05$. We train with learning rate $41e-$ 4 and batch size $8$. When training the Gaussian Head, the Geometry and Image autoencoders are frozen.
@@
To enhance rendering performance and reduce the memory footprint of assets for the viewer, we add an L1 regularizer term to encourage completely transparent Gaussians when they are not necessary, similar to LongLRM [79]. Gaussians with low opacity are then culled before saving the asset. To further reduce file sizes, the model data is quantized in chunks of 256 gaussians (https://github.com/playcanvas/splat-transform).
@@
We train our diffusion model [18] using $v$-parameterization [43] with $T$=1000 timesteps. At inference time, we use DDIM [47] to speed up inference to 50 steps using the same noise schedule as CAT3D [12] except with zero terminal SNR [29].
@@
9 Limitations

### v2

@@ 5.4 Image VAEs generalize poorly to geometry
[…]
@@
We use metrics commonly used for depth estimation [51, 80]—relative error (AbsRel) between the target z-component of the pointmap $z$ and prediction $\hat{z}$, Rel. $\left\lVert z-\hat{z}\right\rVert/z$, and the prediction threshold accuracy, $\delta_{\epsilon}=\operatorname{max}(\nicefrac{{\hat{z}}}{{z}},\nicefrac{{z}}{{\hat{z}}})<\epsilon$ with $\epsilon=1.01$ Additionally, we measure the re-projection error: the mean Euclidean distance on the image plane when re-projecting the point back to the source camera $\Delta uv=\frac{1}{N}\sum_{i}^{N}\left\lVert(u_{i},v_{i})-(\hat{u}_{i},\hat{v}_{i})\right\rVert_{2}$, where $u$ and $v$ are the $x$ and $y$ coordinates, respectively, in pixels on the image. We measure performance separately for synthetic and real data, both at $512\times 512$ resolution.
@@
Our geometry data is not metric, so the scale of the scene is arbitrary. Properly normalizing the data before inputting it to the VAE can significantly impact performance. We consider 3 methods for scaling data. We experiment with (1) scaling the scene such that mean scene depth is $1$, (2) scaling the scene such that the maximum coordinate value in any direction is $1$ and (3) applying a nonlinear contraction function (the sigmoid function) before inputting it to the encoder, and applying its inverse after decoding. Our autoencoder uses (1).
@@ 6 Conclusion
We have presented Bolt3D, a fast feed-forward method that generates detailed 3D scenes in less than 7 seconds. To enable this capability, we propose a 3D scene representation that allows for denoising high-resolution 3D scenes using powerful 2D latent diffusion architectures. We create a large-scale 3D scene dataset to train Bolt3D, and demonstrate that it accurately models ambiguity, enabling high-quality 1-view reconstructions where regression-based methods fail. Our proposed feed-forward approach is capable of reconstructing a wide variety of scenes and reduces the cost of 3D generation by $300\times$ compared to existing optimization-based methods, opening up opportunities for 3D content creation at scale.
@@ Acknowledgments
We would like to express our deepest gratitude to Ben Poole for helpful suggestions, guidance, and contributions. We also thank George Kopanas, Sander Dieleman, Matthew Burruss, Matthew Levine, Daniel Duckworth, Peter Hedman, Songyou Peng, Rundi Wu, Alex Trevithick, Hadi Alzayer, David Charatan, Jiapeng Tang and Akshay Krishnan for valuable discussions and insights. Finally, we extend our gratitude to Shlomi Fruchter, Kevin Murphy, Mohammad Babaeizadeh, Han Zhang and Amir Hertz for training the base text-to-image latent diffusion model. This work was done while S. Szymanowicz was on an internship at Google. He is supported by the EPSRC Doctoral Training Partnerships Scholarship (DTP) EP/R513295/1 and the Oxford-Ashton Scholarship.
@@ 7.1 Videos and interactive results
7.2 Comparison to video models
Video models have emerged as a powerful tool for novel view systhesis. However, using them to reconstruct a 3D asset from a generated video requires distillation [39], similarly to CAT3D, which significantly increases the runtime of such approaches when applied to 3D reconstruction. Nonetheless, we evaluate the quality of novel views generated by one representative approach, MVSplat360 [11]. We evaluate Bolt3D and MVSplat360 on 2-view and 4-view reconstruction on scenes from DL3DV at $512\times 512$ resolution. We take care to input appropriately-sized renders to MVSplat360’s diffusion model to match its training resolution, and crop images appropriately for evaluation. In Tab. 5 we observe that renders from Bolt3D’s 3D scenes are more accurate than the novel views generated by MVSplat360. In Fig. 8 we find that the conditioning often needed for video models (e.g., ViewCrafter [96], MVSplat360, ReconX [39]) can be brittle, resulting in poor accuracy of generated views. In addition, video models are also slow (5.8 minutes for 56 frames) while Bolt3D reconstructs a 3D asset in 7 seconds and renders in real-time.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
2-view MVSplat360 13.97 0.400 0.575 104.83
DL3DV Ours 17.75 0.550 0.392 64.53
4-view MVSplat360 15.42 0.422 0.507 76.91
DL3DV Ours 20.64 0.652 0.311 48.28
[…]
Figure 8: MVSplat360 uses MVSplat for conditioning, which works poorly when there is little or no input view overlap.
@@ 7.3 Ablations
Next, we evaluate performance at a lower resolution, closer to the training setup of baseline methods. In Tab. 9 we illustrate that at a lower $256\times 256$ resolution, our model still outperforms Depthsplat. Finally, we give Depthsplat an advantage by evaluating its performance when receiving wide field-of-view (fov) $256\times 448$ images. In this setting, the wider field-of-view provides more scene coverage by a factor of $\times 1.75$ and more cues for matching across different views. Only in this setting do we find that Depthsplat achieves similar performance to our method.
Rel. $\downarrow$ $\delta_{1.01}$ $\uparrow$ $\Delta uv(\text{px})\downarrow$
@@
1-view RE10K Flash3D $\times256$ 17.70 0.616 0.393
Ours $\times256$ 21.62 0.804 0.202
3-view RE10K Depthsplat $\times256$ 24.69 0.873 0.126
Ours $\times256$ 27.39 0.916 0.103
[…]
Depthsplat $\times256$ 16.16 0.467 0.388
Ours $\times256$ 18.01 0.556 0.320
[…]
Depthsplat $\times256$ 19.64 0.633 0.254
Ours $\times256$ 21.16 0.695 0.231
[…]
Depthsplat $\times256$ 20.64 0.680 0.225
Ours $\times256$ 22.18 0.733 0.206
Table 9: Low-resolution $\times256$ comparisons. Our method outperforms competitors at low resolution when receiving the same input information. Only when DepthSplat [87] receives 1.75 times more input information than our method by ingesting wide-fov images, does its performance become similar to ours.
@@ 8.1 XYZ normalization
The supervising (pseudo-ground truth) data used to train our diffusion model is reconstructed using off-the-shelf 3D reconstruction algorithms (MASt3R) [33]. We transform this 3D reconstruction to the view-space of the first camera, such that all point coordinates and all cameras are relative to this coordinate frame: $\mathbf{\Pi},\mathbf{\mathrm{I}}^{\mathit{xyz}}:=\mathbf{\Pi}\mathbf{\Pi}_{0}^{-1},\mathbf{\mathrm{I}}^{\mathit{xyz}}\mathbf{\Pi}_{0}^{-1}$, where $\mathbf{\Pi}_{0}$ denotes the camera-to-world rigid body transform of the first camera.
@@
We normalize the 3D scale of the reconstructed scenes by applying a per-scene scaling factor $\alpha$ to the camera poses and point coordinates: $\mathbf{\Pi},\mathbf{\mathrm{I}}^{\mathit{xyz}}=\alpha\mathbf{\Pi},\alpha\mathbf{\mathrm{I}}^{\mathit{xyz}}$. This scale factor is chosen such that the mean depth value from the first camera is the same across every scene in our dataset: $\alpha\bar{\mathbf{\mathrm{I}}}^{\mathit{xyz}}_{0}[z]=1$.
@@
where $\left[R~|~T\right]_{\text{w2c}}$ is the world-to-camera transformation matrix. Because scenes are scaled such that the mean depth to the first camera is 1, we can think of $0,0,1]^{\top}$ as the look-at point or center of the scene, so $d=\left\lVert\mathbf{x}_{\text{local}}-[0,0,1]^{\top}\right\rVert$ is the distance of the point to the center of the scene of the local camera. Thus, we compute:
@@
We use a U-Net with full 3D attention on all feature maps up to $32\times 32$, as in CAT3D [18]. Unlike CAT3D, our diffusion model is trained to model the joint distribution of latent appearance and geometry. To this end, we increase the number of channels in the input and output layers of CAT3D’s architecture by 8 to additionally accept geometry latents. The input to our network thus has 8-dimensions for the geometry latents, 8-dimensions for the image latent, 6-dimensions for the camera pose raymaps, and a 1-dimensional mask indicating which views are given as conditioning, yielding an input dimension of $64\!\times\!64\!\times\!23$. We train with the same optimization hyperparameters as [18], except we additionally finetune on $16$ input views with a lower learning rate of 1e51e-5.
@@
We use a pre-trained and frozen image autoencoder similar to that of Stable Diffusion [55]. The geometry encoder has the same architecture, except we increase the channel dimension to additionally accept a 6-dimensional camera pose representation. The decoder is a transformer-based network. We patchify the $64\times 64\times 8$ latent with patch size $2$, thus using a token length $1024$. We use the ViT-B architecture hyperparameters: 12 layers with channel size $768$, with the fully-connected layer consisting of 2 dense layers with GeLU activation function and a hidden MLP dimension $3072$. The linear projection head projects each token to a $16\times 16$ patch. We optimize the parameters of the Autoencoder with the Adam [15] optimizer using constant learning rate 1e41e-4, batch size $512$, Adam parameters $(\beta_{1},\beta_{2})=(0.0,0.99)$. We first train for 3M iterations at $256\times 256$ resolution, followed by fine-tuning at $512\times 512$ for 250k iterations. We use loss weight parameters $\lambda_{1}=3e-9$ and $\lambda_{2}=0.033$.
Foreground masking for synthetic data.
In synthetic data, we apply loss $\mathcal{L}_{rec}$ only on foreground pixels and train the model to additionally output a foreground alpha mask, supervised with binary cross-entropy loss.
@@
We detail the Gaussian head architecture in Fig. 9. The Gaussian Head receives as input the 3-dimensional image, 3-dimensional pointmap, and 6-dimensional raymap encoding the camera pose. Each view is first passed through 3 residual convolutional blocks with the Swish activation function [50] and channel size $128$, followed by $4\times 4$ patchification to token dimension $128$ and full cross attention. We use 3 transformer layers with hidden dimension $128$ and $8$ heads, MLP dimension $512$. The tokens are then unpatchified to original resolution and $128$ channel dimension. Following that, there are another 3 residual convolutional blocks, and a final, unactivated $3\times 3$ convolutional layer that outputs channel size $11$ (33 for color, $3$ for size, $4$ for rotation and $1$ for opacity). The outputs are activated with an exponential function for scale and a sigmoid function for opacity. To facilitate accurate scale prediction, the size output by the network is then multiplied by the z-distance of the gaussian from the camera. The means of Gaussians are not predicted by the Gaussian head, as they are already available from the VAE. The Gaussian head is trained with $8$ input views, leading to a sequence length $131k$. We manage this computational complexity by using FlashAttention [13, 12] and rematerializing gradients on the dot product operation. For losses, we use L2 photometric loss with weight $\lambda=1$ and LPIPS loss weight $\lambda_{LPIPS}=0.05$. We train with learning rate 1e41e-4 and batch size $8$. When training the Gaussian Head, the Geometry and Image autoencoders are frozen.
@@
To enhance rendering performance and reduce the memory footprint of assets for the viewer on our project website, we add an L1 regularizer term to encourage completely transparent Gaussians when they are not necessary, similar to LongLRM [107]. Gaussians with low opacity are then culled before saving the asset. To further reduce file sizes, the model data is quantized in chunks of 256 gaussians (https://github.com/playcanvas/splat-transform).
@@
We train our diffusion model [26] using $v$-parameterization [57] with TT=1000 timesteps. At inference time, we use DDIM [65] to speed up inference to 50 steps using the same noise schedule as CAT3D [18] except with zero terminal SNR [37].
@@
9 Limitations, discussion and future work
Limitations.
@@ 9 Limitations, discussion and future work
Discussion and future work.
To the best of our knowledge, Bolt3D is the first work to explore the architecture and training recipe of a Geometry VAE, and there remain several design choices to be explored. In particular, we chose compress pointmaps over depth due to their resounding success in multi-view reconstruction (Dust3r [80], Mast3r [33]), but concurrent work shows that inferring depth can be complimentary [30] or even advantageous [78]. We leave exploration of these findings in context of 3D generation to future works.
Next, despite making a significant step in feed-forward 3D generation, the quality of 3D generation, Bolt3D lacks in quality compared to optimization-based methods such as CAT3D. We hypothesize this is due to CAT3D generating much more views ($\approx$800, compared to our 16), resulting in more complete scene coverage. Generating more views, perhaps through an anchoring strategy [58], could improve the quality of results, though it would result in a large number of 3D Gaussians.
Finally, Bolt3D generates exclusively static scenes. Perhaps future work could combine multi-view video diffusion models [85] with Bolt3D’s direct geometry generation to generate dynamic 3D scenes in a feed-forward manner.
10 Experimental details
DL3DV scenes.
We ran evaluation on the intersection of our test set and the public test benchmark. In 2-view, 4-view and 6-view setups we used center-crops of views $(0,49)$, $(0,19,29,49)$ and $(0,9,19,29,39,49)$, respectively. The scenes used for evaluation were:
  • 0569e83fdc248a51fc0ab082ce5e2baff15755c53c207f545e6d02d91f01d166,
  • 073f5a9b983ced6fb28b23051260558b165f328a16b2d33fe20585b7ee4ad561,
  • 183dd248f6a86e07c5adf9de8ee2d0abe45b1216331c03678e89634c2e9b1c7f,
  • 1ba74c22670ad047981441581d00f26f4a148d1010bcac7468c615adf5fa4d5d,
  • 389a460ca1995e0658e85fe8e6b520b4e88b370cd6710dfe728b1564bba31aee,
  • 493816813d2d6d248eb3c2b0b77b63e54235266e9a06e270fd0d282f13960493,
  • 50c46cf8b8b22c8d2ffdef8964b05ddbceaef312c9a9ff331d1ecebfd223f72a,
  • 4ae797d07b6d1644c9db6919c8cc8c0d28d72be45108ac7a3abf8dc21b599d83,
  • 565553aa894be621e8b4773cac288e60ad0c2cf7edb621be62b348c9a0f78380,
  • 599ca3e04cae3ec83affc426af7d0d7ab36eb91cd8e539edbc13070a4d455792,
  • 5c8dafad7d782c76ffad8c14e9e1244ce2b83aa12324c54a3cc10176964acf04,
  • 63798f5c6fbfcb4eb686268248b8ecbc8d87d920b2bcce967eeaedfd3b3b6d82,
  • 946f49be73928469000baa5ca04d2573137c5ee6a66362bcf8d130354dca8924,
  • 9e9a89ae6fed06d6e2f4749b4b0059f35ca97f848cedc4a14345999e746f7884,
  • cd9c981eeb4a9091547af19181b382698e9d9eee0a838c7c9783a8a268af6aee,
  • d4fbeba0168af8fddb2fc695881787aedcd62f477c7dcec9ebca7b8594bbd95b.